a bag contain 2 white ball and 3 black balls obtain probability distribution of the of white balls if 2 balls are selected?
Answer by rothauserc(4718)

I assume the balls are selected without replacement
:
let w stand for a white ball and b stand for a black ball, then we have the following outcomes
:
(w,w), (w,b), (b,w), (b,b)
:
Probability(P) of (w,w) is (2/5)*(1/4) = 2/20 = 1/10
:
P of (w,b) = (2/5)*(3/4) = 6/20 = 3/10
:
P of (b,w) = (3/5)*(2/4) = 6/20 = 3/10
:
P of (b,b) = (3/5)*(2/4) = 6/20 = 3/10
:
Probability distribution of white balls
:
Number of White balls|Probability
0 0.3
1 0.6
2 0.1
: